OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form Answer: y + 8 = 3(x + 5) y – 5 = 3(x – 2) y + 5 = 3(x + 2) y – 2 = 3(x – 5)

OpenStudy (chaise):

y=mx+b 5=3x2+b 5=6+b b = -1 y=mx+b y=3x-1

OpenStudy (chaise):

Option two.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

y – 5 = 3(x – 2)